Danielle recently moved to the United States with $10,000 of acceptable currency that had never been in the system before. Assume Danielle deposits the money into First National Bank. If the central bank has set a required reserve ratio of 20 percent, what is the maximum amount of money First National Bank can create?

$50,000

$40,000

$5,000

$4,000

$800
Answer #1

Total money supply or maximum money supply that a bank create = reservemoney multiplier

where money multiplier = 1/r

r = required reserve ratio

Reserve = 10,000

r = 20% = 1/5

  Total money supply or maximum money supply that a bank create = reservemoney multiplier

= reserve(1/r)

= 10,000(1/20%)

= 10,0005

= 50,000

so first option is correct.